# Activation thresholds and expressiveness of polynomial neural networks

### v1

@@ Abstract
Polynomial neural networks have been implemented in a range of applications and present an advantageous framework for theoretical machine learning. A polynomial neural network of fixed architecture and activation degree gives an algebraic map from the network’s weights to a set of polynomials. The image of this map is the space of functions representable by the network. Its Zariski closure is an affine variety known as a neurovariety. The dimension of a polynomial neural network’s neurovariety provides a measure of its expressivity. In this work, we introduce the notion of the activation threshold of a network architecture which expresses when the dimension of a neurovariety achieves its theoretical maximum. In addition, we prove expressiveness results for polynomial neural networks with equi-width architectures.
@@ Introduction
Polynomial neural networks are important in applications and theoretical machine learning. The function spaces and dimensions of neurovarieties for deep linear networks have been studied, and new developments in the polynomial neural network setting have appeared. In particular, results on the choice of the activation degree and the dimension of the neurovariety have improved our understanding of the optimization process of these neural networks and the ability of shallow and deep neural networks to replicate target functions [21, 27].
@@
These theoretical results possess relevant implications. For appropriate datasets, polynomial activation functions can reduce model complexity and computational costs by introducing higher-order interactions between inputs, making it possible to model non-linear phenomena more efficiently. Moreover, polynomial neural networks have been found to perform well in practice in high-impact fields such as healthcare and finance. Polynomial neural networks have been successfully used to predict epidemic behavior [13], quantify the natural frequency of materials [11], analyze financial time series data [14, 25], and improve three-dimensional shape representation for computer vision [32].
@@
The expressiveness of artificial neural networks of different structures and activation has been studied from several perspectives [23, 24, 29]. Substantial literature has shown the expressive power of deep networks, from the perspective of number of linear regions [24], universal approximation and VC dimensions [31, 29], a measure of complexity called “trajectory length” [23], and the exact class of functions representable by ReLU networks of various depths [18]. In another direction, [27] studied the effect of width of convolutional neural networks on the loss landscape, and [22, 17] provided bounds on network width for universal approximation for ReLU networks, while [19] provided such a bound for arbitrary activation functions. The expressiveness of neural networks has also been studied via their topological complexity, for example [4, 16]. In particular, [28] showed that networks with non-increasing architecture (each layer is no wider than the previous) result in a connected decision region. ReLU expressivity results have also been studied using algebraic geometry, specifically tropical geometry [6, 16, 18].
@@
An intuitive way of characterizing the expressiveness of a neural network is by the dimension of the space of representable functions as one allows its parameters to vary. For sigmoid activation functions this has been done in [1], for ReLU activation and non-increasing architectures (as well as a few others) this was done by [7, 16]. For polynomial activation functions, [20, 21] studied this quantity and raised various conjectures, one of which we will resolve in the current paper.
@@
This paper is structured as follows. In Section 1, we review key definitions for polynomial neural networks and introduce the notion of activation thresholds (8). Section 2 determines the expressivity of polynomial neural networks with high activation degree and proves the existence of the activation threshold. The main result resolves a conjecture by Kileel, Trager, and Bruna [20, Conjecture 16] and Kubjas, Li, Wiesmann [21, Conjecture 5.2]. In Section 3, we prove the expressivity of networks with equi-width architecture by computing the activation threshold— in particular, we remove any hypothesis involving “sufficiently high learning degree” by showing the activation threshold of these networks is one. In Section 4 we provide an outlook for the future.
@@ 1 Polynomial neural networks and neurovarieties
For consistency, the notation introduced in this section is intended to agree with [20] and [21] as much as possible.
@@
The architecture of the neural network $F_{\boldsymbol{\theta}}$ is the sequence $\mathbf{d}=(d_{0},\dotsc,d_{L})$. Here, the affine-linear maps are given by
$f_{i}(x):\mathbb{R}^{d_{i-1}}\to\mathbb{R}^{d_{i}},\quad x\mapsto W_{i}x+b_{i},$
where the weights of the neural network are the matrices $W_{1},\dots,W_{L}$ and the biases are the vectors $b_{1},\dotsc,b_{L}$. The parameter set ${\boldsymbol{\theta}}$ consists of these weight matrices and biases. The activation map $\sigma_{i}:\mathbb{R}^{d_{i}}\to\mathbb{R}^{d_{i}}$ is given coordinate-wise by the activation function.
Deep learning consists of approximating a target function $F:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$ by a neural network $F_{\boldsymbol{\theta}}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$ of a chosen architecture $\mathbf{d}=(d_{0},\dotsc,d_{L})$ and activation function. That is, deep learning concerns recovering parameters ${\boldsymbol{\theta}}$ for which $F_{\boldsymbol{\theta}}$ most closely resembles $F$. The map $F_{\boldsymbol{\theta}}$ is associated to the tuple of parameters ${\boldsymbol{\theta}}$ by the parameter map
$\Psi_{\mathbf{d},\sigma}:\mathbb{R}^{N}\to\text{Fun}(\mathbb{R}^{d_{0}},\mathbb{R}^{d_{L}}),\quad{\boldsymbol{\theta}}\mapsto F_{\boldsymbol{\theta}}.$ (1)
In this article, we are interested when the activation function consists of polynomials. The most important case are pure powers, and this leads to the following definition.
@@
$p_{\mathbf{w}}(x)=(W_{L}\circ\sigma_{L-1}\circ W_{L-1}\circ\sigma_{L-2}\circ\cdots\circ\sigma_{1}\circ W_{1})(x)$
@@
In the definition of polynomial neural network, the biases are taken to be zero. The affine-linear map $f_{i}$ in the $i$-th layer is then a linear map that we identify with the matrix $W_{i}$.
An $L$-layer polynomial neural network $p_{\mathbf{w}}$ with architecture $\mathbf{d}=(d_{0},\dotsc,d_{L})$ and activation degree $r$ is a tuple of homogeneous polynomials. The parameter map now takes the matrix tuple $(W_{1},W_{2},\dots,W_{L})$ to a tuple of degree $r^{L-1}$ homogeneous polynomials in $d_{0}$ variables of length $d_{L}$
$\Psi_{\mathbf{d},r}:\mathbb{R}^{d_{1}\times d_{0}}\times\cdots\times\mathbb{R}^{d_{L}\times d_{L-1}}\to(\text{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}}))^{d_{L}},\quad\mathbf{w}\mapsto p_{\mathbf{w}}=\begin{pmatrix}p_{\mathbf{w}_{1}}\\ \vdots\\ p_{\mathbf{w}_{d_{L}}}\end{pmatrix}.$ (2)
To specify an element in the image of $\Psi_{\mathbf{d},r}$, it suffices to identify its vector of coefficients in $\mathbb{R}^{d_{L}\binom{r^{L-1}+d_{0}-1}{d_{0}-1}}\simeq(\text{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}}))^{d_{L}}$.
Remark 3 (Special cases).
Note that if $d_{L}=1$, the co-domain of the parameter map is the space of degree $r^{L-1}$ homogeneous polynomials. If $L=1$ then the ambient space consists of linear maps from $\mathbb{R}^{d_{0}}$ to $\mathbb{R}^{d_{L}}$.
The neuromanifold or functional space associated to $p_{\mathbf{w}}$ is the family of all polynomial neural networks with fixed activation degree $r$ and architecture $\mathbf{d}=(d_{0},d_{1},\dots,d_{L})$ where the weights are allowed to vary.
Definition 4 (Neurovariety).
The image of $\Psi_{\mathbf{d},r}$ is the neuromanifold $\mathcal{M}_{\mathbf{d},r}$, which is a semialgebraic set in $(\text{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}}))^{d_{L}}$. The neurovariety $\mathcal{V}_{\mathbf{d},r}$ is the Zariski closure of $\mathcal{M}_{\mathbf{d},r}$, and is an affine variety in $(\text{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}}))^{d_{L}}$.
The dimension of the neurovariety $\mathcal{V}_{\mathbf{d},r}$ provides a measure of the expressivity of a neural network because it quantifies the degrees of freedom of the space of functions the network produces [20]. In practice, this is the dimension of the neuromanifold $\mathcal{M}_{\mathbf{d},r}$; as $\mathcal{M}_{\mathbf{d},r}$ is semialgebraic, its dimension is the same as that of $\mathcal{V}_{\mathbf{d},r}$.
The size of the weight matrices is a primary factor in determining the dimension of the neurovariety. However, for all network architectures there exists a symmetry in the weight matrices, known as multi-homogeneity. It follows that the dimension of a fiber of the parameter map $\Psi_{\mathbf{d},r}$ is at least $\sum_{i=1}^{L-1}d_{i}$. ([20], Lemma 13).
@@
For all invertible diagonal matrices $D_{i}\in\mathbb{R}^{d_{i}\times d_{i}}$ and permutation matrices $P_{i}\in\mathbb{R}^{d_{i}\times d_{i}}$ ($i=1,\dots,L-1$), the parameter map $\Psi_{\mathbf{d},r}$ returns the same map under the replacement
@@
where $T$ indicates matrix transpose. Thus the dimension of a generic pre-image of $\Psi_{\mathbf{d},r}$ is at least $\sum_{i=1}^{L-1}d_{i}$.
@@
$\text{edim}\mathcal{V}_{\mathbf{d},r}:=\min\left\{d_{L}+\sum_{i=0}^{L-1}(d_{i}d_{i+1}-d_{i+1}),\ d_{L}\binom{d_{0}+r^{L-1}-1}{r^{L-1}}\right\}.$
This definition was introduced in [21] and in [20] it was shown that $\text{edim}\mathcal{V}_{\mathbf{d},r}$ is an upper bound on the dimension of the neurovariety. However, this expected dimension is not always equal to the dimension, as demonstrated by the following example.
@@
The dimension of a neurovareity is subtle even in simple settings. For instance, when the architecture is $(d_{0},d_{1},1)$, a reinterpretation [20, Theorem 9] of the Alexander–Hirschowitz Theorem [2] determines which neurovarieties have the expected dimension. In [21], the authors coin the term defect for the difference between the dimension of a neurovariety from its expected dimension. Examples of $\mathcal{V}_{\mathbf{d},r}$ with defect are found in [21, Table 1] and in [20] as well.
Our main result proves a lower bound on the dimension of a neurovariety for sufficiently high activation degree by resolving a conjecture in [20]. We introduce the notion of the activation threshold, whose existence will follow from our main result Theorem 12.
Problem 8 (Activation threshold).

### v3

# Activation degree thresholds and expressiveness of polynomial neural networks

@@ Abstract
We study the expressive power of deep polynomial neural networks through the geometry of their neurovariety. We introduce the notion of the activation degree threshold of a network architecture to express when the dimension of the neurovariety achieves its theoretical maximum. We prove the existence of the activation degree threshold for all polynomial neural networks without width-one bottlenecks and demonstrate a universal upper bound that is quadratic in the width of largest size. In doing so, we prove the high activation degree conjecture of Kileel, Trager, and Bruna. Certain structured architectures have exceptional activation degree thresholds, making them especially expressive in the sense of their neurovariety dimension. In this direction, we prove that polynomial neural networks with equi-width architectures are maximally expressive by showing their activation degree threshold is one.
@@ Introduction
Characterizing the functional space of artificial neural networks is a fundamental project in deep learning theory. When the functional space is semialgebraic, its geometric properties offer insight and mathematical intuition into the learning dynamics. Such a characterization is a two-way street: it is foundational to explaining empirical phenomena and offers the possibility of designing innovative architectures based on desirable theoretical and practical properties encoded in the space. Semialgebraic machine learning models form a rich class that encompasses, for example, Rectified Linear Unit (ReLU) and polynomial neural networks.
The study of deep polynomial neural networks offers a distinct advantage in theoretical machine learning because the functional spaces of such networks are described by algebraic varieties. In particular, a polynomial neural network of fixed architecture and activation degree gives an algebraic map from the network’s weights to a set of polynomials. The image of this map is the space of functions that can be represented by the network. Its closure is an affine variety known as a neurovariety. The algebro-geometric invariants of a neurovariety, such as its dimension, degree, and singularities, encode information about the network’s training dynamics and expressiveness. For a detailed discussion on neuroalgebraic geometry detailing the relationship between invariants of a neurovariety and certain fundamental aspects of machine learning, see [30].
@@
The most basic instantiation of a polynomial neural network occurs when the activation function is linear. Linear networks have received wide attention as simplified models for understanding the learning behavior of deep neural networks [23, 6]. Even in the case of linear networks, when the neuromanifold is a determinantal variety, there is interesting geometry to explore [42, 26, 27]. New developments in the general polynomial neural network setting have demonstrated the cross-disciplinary connections between algebraic geometry and machine learning and offered results on the viability of polynomial neural networks for applications [25, 40]. Results on the choice of the activation degree and the dimension of the neurovariety have improved our understanding of the optimization process of these neural networks and the ability of shallow and deep neural networks to replicate target functions [28, 35, 5].
While polynomial neural networks with fixed activation degree are not universal approximators [21], they become so when the activation degree is allowed to vary. An alternative measure of a polynomial neural network’s expressiveness is the dimension of its neurovariety, which can be defined as the maximum number of independent vectors in the tangent space at a generic point. From the perspective of algebraic geometry, the dimension of an algebraic variety is perhaps its most basic invariant. Characterizations of the neurovariety associated with a neural network represent a distinct description of a network’s expressiveness from universal approximation theorems. The dimension of the neurovariety is a simple and precise measure of the degrees of freedom of the model that admits exact computation. In contrast to universal approximation theorems, the dimension of the neurovariety is a measure of the expressiveness of a given polynomial neural network. Thus, studying polynomial network expressiveness through the dimension of the neurovariety complements universal approximation results. For an explicit toy example illustrating neurovariety dimension, see [30, Section A].
@@
In practice, neural networks with polynomial output have been found to perform well in high-impact fields such as healthcare and finance, especially where the nature of the data is polynomial. Such applications include the prediction of epidemic behavior [15], quantification of the natural frequency of materials [13], analysis of financial time series data [16, 33] and the improvement of three-dimensional shape representation for computer vision [44]. In these applications, by introducing higher-order interactions between inputs, polynomials model non-linear phenomena more efficiently.
Other applications [22] take advantage of the fact that a homogeneous polynomial of degree $d$ in $n$ variables can be identified with a symmetric tensor in $\left(\mathbb{R}^{n}\right)^{\otimes d}$ that collects its coefficients. This description is also useful for obtaining theoretical results. It is exhaustive for shallow networks, where the function space coincides with a set of symmetric tensors of bounded rank [25]. Moreover, it provides a lens for studying the loss surface of neural networks from an optimization perspective [43, 5].
@@
The expressiveness of artificial neural networks of different structures and activation has been studied from several perspectives [31, 32, 37]. Substantial literature has shown the expressive power of deep networks, from the perspective of number of linear regions [32], universal approximation and VC dimensions [41, 37], a measure of complexity called “trajectory length” [31], and the exact class of functions representable by ReLU networks of various depths [20]. In another direction, [35] studied the effect of width of convolutional neural networks on the loss landscape, and [29, 19] provided bounds on network width for universal approximation for ReLU networks, while [24] provided such a bound for arbitrary activation functions. The expressiveness of neural networks has also been studied via their topological complexity, for example [7, 18]. In particular, [36] showed that networks with non-increasing architecture result in a connected decision region. ReLU expressiveness results have also been studied using algebraic geometry, specifically tropical geometry [9, 18, 20].
@@
An intuitive way of characterizing the expressiveness of a neural network is by the dimension of the space of representable functions as one allows its parameters to vary. For sigmoid activation functions this has been done in [1], for ReLU activation and non-increasing architectures, as well as a few others, this was done by [10, 18]. For polynomial activation functions, [25, 28] studied this quantity and raised various conjectures, one of which we will resolve in the present paper.
@@
This paper is structured as follows. In Section 1, we review key definitions for polynomial neural networks and introduce the notion of activation degree thresholds (Definition 10). In Section 2, we prove the existence of activation degree thresholds and determine the fibers of the parameter map of polynomial neural networks with high activation degree, resolving [25, Conjecture 16] and [28, Conjecture 5.2], raised by [14]. In Section 3, we prove the expressiveness of networks with equi-width architecture by computing the activation degree threshold— in particular, we remove any hypothesis involving “sufficiently high learning degree” by showing the activation degree threshold of these networks is one. In Section 4 we provide an outlook for the future, focusing on research directions opened by activation degree thresholds.
@@ 1 Polynomial neural networks and neurovarieties
1.1 Background
@@
The architecture of the neural network $F_{\boldsymbol{\theta}}$ is the sequence $\mathbf{d}=(d_{0},d_{1},\dotsc,d_{L})$. Here, the affine-linear maps are given by
$f_{i}:\mathbb{R}^{d_{i-1}}\to\mathbb{R}^{d_{i}},\quad x\mapsto W_{i}x+b_{i},$
where the weights of the neural network are the matrices $W_{1},W_{2},\dots,W_{L}$ and the biases are the vectors $b_{1},\dotsc,b_{L}$. The parameter set ${\boldsymbol{\theta}}$ consists of these weight matrices and biases. The activation map $\sigma_{i}:\mathbb{R}^{d_{i}}\to\mathbb{R}^{d_{i}}$ is given coordinate-wise by the activation function.
Deep learning consists of approximating a target function $F:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$ by a neural network $F_{\boldsymbol{\theta}}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$ of a chosen architecture $\mathbf{d}=(d_{0},d_{1},\dotsc,d_{L})$ and activation function. That is, deep learning concerns recovering parameters ${\boldsymbol{\theta}}$ for which $F_{\boldsymbol{\theta}}$ most closely resembles $F$. If $\mathbb{R}^{N}$ is the ambient space of the parameter set, the map $F_{\boldsymbol{\theta}}$ is associated to the $N$-tuple of parameters ${\boldsymbol{\theta}}$ by the parameter map
$\Psi_{\mathbf{d},\sigma}:\mathbb{R}^{N}\to\operatorname{Fun}(\mathbb{R}^{d_{0}},\mathbb{R}^{d_{L}}),\quad{\boldsymbol{\theta}}\mapsto F_{\boldsymbol{\theta}}.$ (1)
In this article, we are interested in the setting in which the activation function is polynomial. The most important cases are when the activation function is a pure power, and this leads to the following definition.
@@
$p_{\mathbf{w}}(x)=(W_{L}\circ\sigma_{L-1}\circ W_{L-1}\circ\sigma_{L-2}\circ\cdots\circ\sigma_{1}\circ W_{1})(x)$ (2)
@@
In the definition of a polynomial neural network, the biases are taken to be zero. The affine-linear map $f_{i}$ in the $i$-th layer is then a linear map that we identify with the matrix $W_{i}$.
An $L$-layer polynomial neural network $p_{\mathbf{w}}$ with architecture $\mathbf{d}=(d_{0},d_{1},\dotsc,d_{L})$ and activation degree $r$ is represented by a tuple of homogeneous polynomials. Precisely, the parameter map now takes the matrix tuple $(W_{1},W_{2},\dots,W_{L})$ to a tuple of degree $r^{L-1}$ homogeneous polynomials in $d_{0}$ variables of length $d_{L}$
$\begin{split}\Psi_{\mathbf{d},r}:\mathbb{R}^{d_{1}\times d_{0}}\times&\cdots\times\mathbb{R}^{d_{L}\times d_{L-1}}\to(\text{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}}))^{d_{L}},\\ \mathbf{w}&\mapsto p_{\mathbf{w}}=\begin{pmatrix}p_{\mathbf{w}_{1}}\\ \vdots\\ p_{\mathbf{w}_{d_{L}}}\end{pmatrix}.\end{split}$ (3)
To specify an element in the image of $\Psi_{\mathbf{d},r}$, it suffices to identify its vector of coefficients in $\mathbb{R}^{d_{L}\binom{r^{L-1}+d_{0}-1}{d_{0}-1}}\simeq\left(\text{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}})\right)^{d_{L}}$.
Remark 3 (Special case).
Note that if $d_{L}=1$, the codomain of the parameter map (3) is the space of degree $r^{L-1}$ homogeneous polynomials.
The neuromanifold or functional space associated to $p_{\mathbf{w}}$ is the family of all polynomial neural networks with fixed activation degree $r$ and architecture $\mathbf{d}$ where the weights are allowed to vary. The image of $\Psi_{\mathbf{d},r}$ is the neuromanifold $\mathcal{M}_{\mathbf{d},r}$, which is a semialgebraic set in $(\text{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}}))^{d_{L}}$. (In spite of the terminology, the neuromanifold is not in general a smooth manifold, and its singularities can affect the behavior of the corresponding model, see [4, 30, 40].) The neurovariety $\mathcal{V}_{\mathbf{d},r}$ is the Zariski closure of $\mathcal{M}_{\mathbf{d},r}$, and is a real affine variety in $(\text{Sym}_{r^{L-1}}(\mathbb{R}^{d_{0}}))^{d_{L}}$.
The dimension of the neurovariety $\mathcal{V}_{\mathbf{d},r}$ provides a measure of the expressiveness of a neural network because it quantifies the degrees of freedom of the space of functions the network produces [25]. In practice, this is the dimension of the neuromanifold $\mathcal{M}_{\mathbf{d},r}$; as $\mathcal{M}_{\mathbf{d},r}$ is semialgebraic, its dimension is the same as that of $\mathcal{V}_{\mathbf{d},r}$.
For all network architectures there exists a symmetry in the weight matrices, known as multi-homogeneity. It follows that the dimension of a fiber of the parameter map $\Psi_{\mathbf{d},r}$ is at least $\sum_{i=1}^{L-1}d_{i}$ ([25], Lemma 13).
@@
For all invertible diagonal matrices $D_{i}\in\mathbb{R}^{d_{i}\times d_{i}}$ and permutation matrices $P_{i}\in\mathbb{R}^{d_{i}\times d_{i}}$ ($i=1,\dots,L-1$), the parameter map $\Psi_{\mathbf{d},r}$ returns the same neural network under the replacement
@@
where $T$ denotes the matrix transpose. Thus the dimension of a generic pre-image of $\Psi_{\mathbf{d},r}$ is at least $\sum_{i=1}^{L-1}d_{i}$.
@@
$\mathrm{edim}\mathcal{V}_{\mathbf{d},r}:=\min$$\Bigg{\{}d_{L}+\sum_{i=0}^{L-1}(d_{i}d_{i+1}-d_{i+1}),\quad d_{L}\binom{d_{0}+r^{L-1}-1}{r^{L-1}}\Bigg{\}}.$ (4)
This definition was introduced in [28] and in [25] it was shown that $\operatorname{edim}\mathcal{V}_{\mathbf{d},r}$ is an upper bound on the dimension of the neurovariety.
@@
In the case when $L=1$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension $d_{0}d_{1}$, as the map $\Psi_{\mathbf{d},r}$ is an isomorphism. The function space is parameterized by the $d_{1}\times d_{0}$ matrix $W_{1}$.
However, the expected dimension is not always equal to the dimension, as demonstrated by the following example.
Example 7.
The dimension of a neurovariety is subtle even in simple settings. For instance, when the architecture is $(d_{0},d_{1},1)$, a reinterpretation [25, Theorem 9] of the Alexander–Hirschowitz Theorem [2] determines which neurovarieties have the expected dimension. In [28], the authors coin the term defect for the difference between the expected dimension of a neurovariety and its dimension:
$\operatorname{defect}(\mathcal{V}_{\mathbf{d},r}):=\operatorname{edim}(\mathcal{V}_{\mathbf{d},r})-\dim(\mathcal{V}_{\mathbf{d},r}).$
Examples of neurovarieties $\mathcal{V}_{\mathbf{d},r}$ with positive defect are found in [28, Table 1] and [25].
Remark 8.
The notion of neurovariety dimension as a measure of network expressiveness has been utilized for designing neural networks. For example, [22] employs a hybrid network with monomial and LeakyReLU activation functions to determine the entanglement type of quantum states. Because their architecture is shallow, they apply the Alexander-Hirschowitz Theorem when choosing the number of neurons in each layer to guarantee that the associated neurovariety achieves the expected dimension. Such applications motivate the project of describing the neurovariety dimension for broader classes of architectures. Our Theorem 18 and Theorem 22 are examples of this kind of result.
1.2 Main result and example
One of our main results proves a lower bound on the dimension of a neurovariety for sufficiently high activation degree. We introduce the notion of the activation degree threshold, whose existence will follow from our Theorem 18.
Problem 9.